OpenStudy (dmezzullo):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. y = 4x - 11 y = 4x + 11 y = 4x + 7 y = 4x - 7

OpenStudy (dmezzullo):

@yely_p

OpenStudy (anonymous):

Point-Slope form. y - y1 = m(x - x1) Fill in the values. y - (-1) = 4(x - (-3)) y + 1 = 4(x + 3) Change to Slope-intercept form. y = mx + b y + 1 = 4(x + 3) Distribute. y + 1 = 4x + 12 Subtract 1. y = 4x + 11